OpenStudy (anonymous):

Simplify: 5c^-9 d^-2 e^3 ________________ 2c^-3 d^5 e^-10

OpenStudy (chihiroasleaf):

have you learned about division rule of exponent? :)

OpenStudy (anonymous):

subtract exponents?

OpenStudy (anonymous):

add the exponts?

OpenStudy (chihiroasleaf):

yes..., \[\Large \frac{x^{a}}{x^b} = x^{a-b} \] \[\Large x^{a} \times {x^b} = x^{a+b} \] now.. can you use these in your question?